DK4L:

The formula for the area of a triangle is A=1/2bh Find the area of a triangle with a base (b) 9 feet and a height (h) 16 feet. The area of the triangle is blank square feet.

nini01:

A=hbb 2=16·9 2=72ft²

XioGonz:

Multiply 9*16=144 Divide 144/2=72

emilysmith1105:

ok so you want to replace the variables (letters) with the values given for them. so it will be A= 1/2(9x16) now you multiply b x h to get 144. so replace that to get A=1/2(144). now simplify- A=72 sq ft.

DK4L:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @emilysmith1105 ok so you want to replace the variables (letters) with the values given for them. so it will be A= 1/2(9x16) now you multiply b x h to get 144. so replace that to get A=1/2(144). now simplify- A=72 sq ft. \(\color{#0cbb34}{\text{End of Quote}}\) ok i got it thnx

emilysmith1105:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @DK4L \(\color{#0cbb34}{\text{Originally Posted by}}\) @emilysmith1105 ok so you want to replace the variables (letters) with the values given for them. so it will be A= 1/2(9x16) now you multiply b x h to get 144. so replace that to get A=1/2(144). now simplify- A=72 sq ft. \(\color{#0cbb34}{\text{End of Quote}}\) ok i got it thnx \(\color{#0cbb34}{\text{End of Quote}}\) np!